Ben hiked 4 4/5 miles on Saturday. He hiked 1 3/4 miles more on Sunday than Saturday. How many more miles did he hike in all?
Answer by stanbon(75887)

Ben hiked 4 4/5 miles on Saturday. He hiked 1 3/4 miles more on Sunday than Saturday. How many more miles did he hike in all?
------------------
Total = 4 4/5 + 4 4/5 + 1 3/4
----
Total = 8 8/5 + 1 3/4
----
Total = 48/5 + 7/4
-----
Total = (4*48 + 5*7)/20
----
Total = 227/20
----
Total = 11 7/20
----------------
Cheers,
Stan H.
---------------